Compute the future value in year 10 of a $1,000 deposit in year 1, and another $1,500 deposit at the end of year 4 using an 8 percent interest rate.
Answer #1
N= 10 − 1 = 9N= 10 − 4 = 6
I= 8I= 8
PV= 1000PV= 1500
PMT= 0PMT= 0
CPT FV= 1999.00CPT FV= 2380.31
1999.00 + 2380.31= 4379.31


answered by: Kraer4342